find solution to 8.4=3/4 log(x1(10^2.4))
Answer by stanbon(75887)

8.4=(3/4)log(x/(10^2.4))
---
log(x) - log(10^2.4) = 11.2
---
log(x) - 2.4 = 11.2
---
log(x) = 13.6
---
x = 10^13.6
==================
Cheers,
Stan H.
----------